OpenStudy (anonymous):

Consider a sphere of radius r. If the radius is multiplied by , how is the volume of the resulting sphere affected?

Directrix (directrix):

If the radius is multiplied by x, then the volume is multiplied by x^3

Directrix (directrix):

Did you leave something out of the problem? If the radius is multiplied by , --> multiplied by what?

OpenStudy (anonymous):

sorry, if the radius is multiplied by 2

Directrix (directrix):

Then the volume is multiplied by 8. All spheres are similar. The cube of the scale factor is the ratio of the volumes (1/2) ^3 = V1/V2 1/8 = V1/V2